# Capturing Turbulence with Numerical Dissipation: a Simple Dynamical Model for Unresolved Turbulence in Hydrodynamic Simulations

### v2

@@ 2.3 Semi-implicit Large-Eddy Simulation (SLES)
In this paper, we propose a model that combines features of both implicit and explicit LES described above. Specifically, we propose using the local numerical dissipation rate—that provides a reasonable approximation for the turbulent cascade rate (see Section 2.1)—as a source term for the explicitly followed subgrid turbulent energy. Such a model can be thought of as an explicit LES model where the source term $\Sigma$ in Equation (1) is computed implicitly analogously to ILES. Conversely, this model can also be thought of as an ILES, where subgrid turbulence and thermal energies are followed separately and are subject to their own explicit sink and source terms. For these reasons, we dub this model Semi-implicit LES (SLES) to highlight the analogy with both explicit and implicit LES.
The advantage of such a model over an explicit LES is that it does not require assuming and implementing closure relations for viscous subgrid terms. It only requires computing the local dissipation rate which can be done straightforwardly as described below. At the same time, unlike ILES, the model explicitly tracks the evolution of subgrid turbulent energy that can be used in the subgrid models for other processes that are sensitive to unresolved turbulence.
@@
When additional non-thermal energy like subgrid $e_{\rm turb}$ is added to the system, one needs to choose how to partition the difference $e_{\rm tot}-e_{\rm kin}$ between $e_{\rm turb}$ and $e_{\rm th}$ (semenov22, see also gupta21 for this problem in the context of cosmic ray hydrodynamics). In particular, the difference $e_{\rm tot}-e_{\rm kin}$ implicitly contains the adiabatic change of $e_{\rm turb}$ and $e_{\rm th}$ as well as the kinetic energy dissipated locally during the hydrodynamic step, and therefore this choice will dictate how this dissipated energy is split between $e_{\rm turb}$ and $e_{\rm th}$. To put all dissipated energy into $e_{\rm turb}$ as envisioned in our simple SLES model, one simply needs to update $e_{\rm turb}$ after the hydrodynamic step as
@@
In principle, one could not follow $e_{\rm turb}$ explicitly but still apply all relevant subgrid terms (except for turbulent viscosity) to other quantities, computing $e_{\rm turb}$ from Equation (2) when needed. However, it is advantageous to follow $e_{\rm turb}$ explicitly so that it can be followed even in the regimes where the right-hand side of Equation (2) is negative. Avoiding such an unphysical behavior requires enforcing local entropy-stability of the numerical scheme (i.e., entropy must only increase) which is generally not guaranteed (see also Section 5). Instead, in our proof-of-concept implementation, we simply turn off this implicit source term in regions with negative dissipation. We find that only a small fraction of cells is affected in our turbulence tests below and the overall results remain the same.
Finally, it is worth noting, that the above model does not affect the adiabatic evolution of gas. Indeed, only the fact that the residual energy ($e_{\rm tot}-e_{\rm kin}$) is partitioned between $e_{\rm th}$ and $e_{\rm turb}$ distinguishes it from ILES, however, given that both energies obey the same $\gamma=5/3$ equation of state, the pressure exerted on the gas remains exactly the same. This is also true when the dissipation term ($\epsilon$ in Equation (1)) is modeled explicitly, as it only changes the partitioning between $e_{\rm th}$ and $e_{\rm turb}$ but not their spatial distribution. This implies that the performance of the code in all idealized adiabatic tests without radiative cooling, like shock tube or point-source explosion, remains exactly the same as without this model. In the presence of cooling, on the other hand, the solution becomes different, as the model helps to retain some of the dissipated kinetic energy in the form of $e_{\rm turb}$ which decays on a timescale different from the thermal cooling timescale.
@@ 2.4 Turbulence decay, diffusion, and extra sources
Other than $\Sigma$, the rest of the explicit source terms in the second line of Equation (1), are the same in LES and SLES. Here we describe the adopted closures for these terms.
@@
The same term enters the equation for thermal energy (which is analogous to Equation (1)) but with the positive sign, implying that the dissipated subgrid turbulence appears as heat which is then subject to radiative losses. The amount of dissipated subgrid turbulence, $\Delta e_{\rm turb}$, during a timestep, $\tau$, is calculated from the analytic solution of $de_{\rm turb}/dt=-\epsilon$:
@@
with $\Delta e_{\rm turb}$ being subtracted from $e_{\rm turb}$ and added to $e_{\rm th}$ after each step.
Next, for the turbulent diffusion term ($\boldsymbol{\nabla}\boldsymbol{\cdot}\kappa\boldsymbol{\nabla}e_{\rm turb}$) we adopt turbulent diffusivity of $\kappa\equiv C^{\prime}_{\kappa}\sigma\,\Delta=C_{\kappa}\sqrt{e_{\rm turb}/\rho}\Delta$ with $C_{\kappa}=\sqrt{2}\,C^{\prime}_{\kappa}=0.4$ (schmidt14; schmidt). A similar term is applied to all advected scalars including thermal energy and chemical abundances when the latter are followed explicitly. In our tests, the effect of explicit subgrid turbulent diffusion is generally smaller than numerical diffusivity except for the regions with very high $\sigma$ such as interiors of supernovae bubbles in galaxy simulations (Section 4). The effect of turbulent diffusion is significantly more important in Lagrangian and quasi-Lagrangian (moving-mesh) codes owing to their reduced advection errors (e.g., rennehan21). As the choice of $C_{\kappa}$ does not affect our results, we do not consider it as a tunable parameter of the model even though we keep the diffusion term on in the tests presented below. Turning off this term completely affects our results only weakly.
Finally, the explicit source term $S$ can be used to source subgrid turbulence via various unresolved processes including stellar feedback. For example, a fraction of supernova energy injected during a timestep can be added directly to $e_{\rm turb}$ via this term. In our galaxy simulations presented in Section 4, we do not perform such an explicit injection (i.e., $S\equiv 0$) but we keep this term in Equation (1) for completeness.
@@ 3 Decaying supersonic turbulence
In this section, we use a decaying supersonic isothermal turbulence test to calibrate and investigate the performance of our new Semi-implicit Large Eddy Simulation (SLES) model. To this end, we use a high-resolution simulation to predict the distribution of turbulent energy on small scales. This distribution is then compared with low-resolution resimulations of the same initial conditions where the small-scale turbulence is predicted using the SLES model. As we will show, despite its simplicity, the SLES model is able to capture multiple non-trivial features of the DNS simulations, such as the time evolution of the mean turbulent energy, its dependence on scale, and the correlation between gas density and turbulent velocity.
@@ 3.1 Numerical setup and analysis
Figure 2: Gas density (gray) and small-scale turbulent energy slices (orange) in the coarse-grained high-resolution simulation (DNS; the top set of panels) and low-resolution resimulations of the same ICs with the SLES model (the bottom set of panels). The maps are shown at $t=2/3\;t_{\rm cross,0}$. The leftmost panel in the top row shows the density slice at the native resolution of the DNS, while the other columns show coarse-grained values, with density being volume-weighted and $e_{\rm turb}$ computed from Equation (8); the corresponding grid size is indicated at the top. The bottom panels show slices at the same timestep from low-resolution resimulations using corresponding grid sizes where $e_{\rm turb}$ is predicted using the SLES model. Thus, in total, four simulations are shown in the plot as indicated by the green shaded background: one DNS and three SLES resimulations. Low-resolution SLES resimulations produce smoother solutions than the DNS but they capture main global features of the small-scale turbulence, in particular, the correlation with gas density.
@@
To initialize the decaying isothermal turbulence test, we use the initial conditions from the code comparison project of kritsuk11. The initial sonic Mach number in this setup is $\mathcal{M}\approx 9$ and it decreases with time as turbulence is allowed to decay. The results below assume units in which the box size is $L_{\rm box}=1$, the average density is $\langle\rho\rangle=1$, and the constant sound speed is $c_{\rm s}=1$.
@@
After the start, the simulation undergoes a brief initial transient stage that lasts approximately one turbulent box-crossing time (see the slight bump in the dashed lines at $t<t_{\rm cross,0}$ in Figures 3 and 4 below). This transient is mainly caused by two factors. First, the initial conditions were created using a higher order method (PPM) than we employ in our run. Second, the original simulation includes magnetic pressure which we removed in our runs. This pressure is however dynamically subdominant (corresponding to the Alfvén Mach number of $\approx 4.4$). Overall, the strength of the transient is rather small and we present our analyses at sufficiently late times, $>t_{\rm cross,0}$, after it is dissipated.
@@
As a reference for calibration and testing of the model, we run a high-resolution version of this setup with a $1024^{3}$ uniform grid. We do not model physical viscosity; instead, kinetic energy is dissipated numerically on scales close to resolution. However, for comparisons with the SLES model, we compute the turbulent energy on scales larger than these numerical dissipation scales, on which the mean turbulent energy is converged ($L\geq L_{\rm box}/64$; see Appendix A). For this reason, we henceforth refer to this reference simulation as a “direct numerical simulation” (DNS) even though, formally, the physical dissipation is absent in this simulation (the so-called “Coarse DNS” approach).
@@
The distribution of small-scale turbulence is derived by coarse-graining the direct simulation and averaging the properties of the flow on the scales below the coarse-graining scale. To this end, we use uniform $N^{3}$ grids, where $N<1024$ and is a power of 2 such that each coarse cell is a cubic volume with $1024/N$ cells on a side. To respect conservation laws, the averages of all conservative quantities are volume-weighted. For example, the coarse-grained density is computed as:
@@
where $dV\equiv(L_{\rm box}/1024)^{3}$ and $dV_{N}\equiv(L_{\rm box}/N)^{3}$ are cell volumes of the original $1024^{3}$ and coarse-grained $N^{3}$ grids.
@@
Implementation-wise, we perform the above averaging by exploiting the nested oct structure of the computational mesh in the ART code. Specifically, we use a static $16^{3}$ root grid where each cell is split into octs until the target uniform resolution of $1024^{3}$ is achieved, which corresponds to 7 refinement levels. Then, starting from the highest refinement level, hydrodynamic quantities are averaged between the cells in a given oct and saved level-by-level providing a hierarchy of coarse-grained versions of the direct simulation.
The top row of Figure 2 shows the original high-resolution density map (the leftmost panel) and the series of coarse-grained versions using uniform grids with $N=128$, 64, and 32 cells on a side at $t=2/3\;t_{\rm cross,0}$, where $t_{\rm cross,0}$ is the initial turbulent box-crossing time. [^4] The specific time at which the maps are shown is picked such that the small-scale distribution of $e_{\rm turb}$ has time to develop but the initial large-scale turbulent energy has not dissipated away. The top (gray) and bottom (orange) subsets of maps show, respectively, coarse-grained density (Equation (6)) and the turbulent energy on the scales below the coarse-graining scale ($e_{\rm turb}$; Equation (8)). The distribution of $e_{\rm turb}$ shows a clear correlation with density: the higher turbulent energy corresponds to denser regions. In addition, the global mean value of $e_{\rm turb}$ increases with decreasing resolution of the coarse-graining grid, i.e., the map becomes more orange from left to right. This is because larger scales contain more turbulent energy.
@@
The results of such low-resolution resimulations with the SLES model are shown in the bottom set of panels. These low-resolution resimulations reproduce the overall features of the direct simulation such as the locations of the high-density structures and the correspondence between high-density and high-$e_{\rm turb}$ regions. The solution is significantly smoother than in the DNS owing to higher numerical viscosity, i.e., larger numerical dissipation scales. The SLES model also captures the increase of the mean $e_{\rm turb}$ at a lower resolution.
Having such a set of simulations, the SLES model can be calibrated and tested by comparing its predictions with the coarse-grained DNS. The cell-by-cell comparison of $e_{\rm turb}$ is complicated due to the strong effect of resolution on the flow structure. Indeed, even minute changes due to, e.g., variations in the time-stepping scheme can cause slight offsets of the large-scale features like the locations of the shock fronts, leading to large cell-to-cell differences between the two runs (see, e.g., Figure 12 in gnedin18). For that reason, in the subsequent sections, we compare global statistics: the mean $e_{\rm turb}$ averaged over the entire box (which we also use to calibrate the model in Section 3.2), the normalization, slope, and scatter of the relation between gas density and $e_{\rm turb}$, and the power spectra of resolved kinetic energy, density, and $e_{\rm turb}$.
@@ 3.2 Model calibration
Figure 3: Comparison of the global mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, in coarse-grained DNS (dashed line) and low-resolution SLES resimulations with different values of the dissipation parameter, $C_{\epsilon}$ (colored lines). For this comparison, we used the grid size of 323, i.e., the cell size $\Delta\equiv L_{\rm box}/32$. In the DNS, $e_{\rm turb}$ is computed following Equation (8) while in the low-resolution resimulations, it is tracked by the SLES model. The mean $E_{\rm turb}$ is normalized to the total initial kinetic energy, i.e., $E_{\rm kin,0}\equiv E_{\rm turb}(<L_{\rm box})$ at $t=0$. The time-axis is normalized by the initial box-crossing time of turbulence, $t_{\rm cross,0}\equiv 0.5\;L_{\rm box}/\sigma_{0}$, or $\approx 0.058$ sound-crossing times for the initial Mach number of $\mathcal{M}\approx 8.6$. The value of $C_{\epsilon}=1.6$ provides the best match to the global $E_{\rm turb}$ and therefore we will use it as our fiducial choice in the rest of the paper. The agreement with the DNS can be further improved by making $C_{\epsilon}$ dependent on the local subgrid Mach number as shown with the thin line and described in the text.
The simplest version of the SLES model involves a single parameter that needs to be selected: $C_{\epsilon}$, describing the rate of subgrid turbulence dissipation in units of local eddy cell-crossing time (Equation (3)). We calibrate the value of $C_{\epsilon}$ by matching the mean subgrid turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$.
@@
Note that as the model contains only a single parameter, only a single instantaneous value of $E_{\rm turb}$ from one of the timesteps could be used for calibration. The fact that the SLES model also reasonably closely reproduces the temporal behavior of $E_{\rm turb}$ is non-trivial and can be viewed as a test of the SLES model rather than a part of calibration (see also Section 3.3.1). Indeed, as the global Mach number of decaying turbulence is decreasing, the values of $E_{\rm turb}$ separated by more than the global turbulence crossing time, $t_{\rm cross,0}$, can be interpreted as independent realizations of different Mach number regimes.
Finally, note that, although the simple SLES model captures the overall behavior of $E_{\rm turb}$ in the DNS, there are some differences in detail. In particular, around $t\sim 2\;t_{\rm cross,0}$, the value of $E_{\rm turb}$ in the SLES model overshoots DNS by $\sim 20\%$ and then decays at a somewhat faster rate resulting in $\sim 20\%$ lower $E_{\rm turb}$ than DNS at $t\gtrsim 8\;t_{\rm cross,0}$.

### v3

@@ 2.3 Semi-implicit Large-Eddy Simulation (SLES)
In this paper, we propose a model that combines features of both implicit and explicit LES described above. Specifically, we propose using the local numerical dissipation rate—which provides a reasonable approximation for the turbulent cascade rate (see Section 2.1)—as a source term for the explicitly followed subgrid turbulent energy. Such a model can be thought of as an explicit LES model where the source term $\Sigma$ in Equation (1) is computed implicitly, analogously to ILES. Conversely, this model can also be thought of as an ILES, where subgrid turbulence and thermal energies are followed separately and are subject to their own explicit sink and source terms. For these reasons, we dub this model Semi-implicit LES (SLES) to highlight the analogy with both explicit and implicit LES.
The advantage of such a model over an explicit LES is that it does not require assuming and implementing closure relations for viscous subgrid terms. It only requires computing the local dissipation rate, which can be done straightforwardly as described below. At the same time, unlike ILES, the model explicitly tracks the evolution of subgrid turbulent energy that can be used in the subgrid models for other processes that are sensitive to unresolved turbulence.
@@
When additional nonthermal energy like subgrid $e_{\rm turb}$ is added to the system, one needs to choose how to partition the difference $e_{\rm tot}-e_{\rm kin}$ between $e_{\rm turb}$ and $e_{\rm th}$ (semenov22, see also gupta21 for this problem in the context of cosmic-ray hydrodynamics). In particular, the difference $e_{\rm tot}-e_{\rm kin}$ implicitly contains the adiabatic change of $e_{\rm turb}$ and $e_{\rm th}$ as well as the kinetic energy dissipated locally during the hydrodynamic step, and therefore, this choice will dictate how this dissipated energy is split between $e_{\rm turb}$ and $e_{\rm th}$. To put all dissipated energy into $e_{\rm turb}$ as envisioned in our simple SLES model, one simply needs to update $e_{\rm turb}$ after the hydrodynamic step as
@@
In principle, one could not follow $e_{\rm turb}$ explicitly but still apply all relevant subgrid terms (except for turbulent viscosity) to other quantities, computing $e_{\rm turb}$ from Equation (2) when needed. However, it is advantageous to follow $e_{\rm turb}$ explicitly so that it can be followed even in regimes where the right-hand side of Equation (2) is negative. Avoiding such an unphysical behavior requires enforcing local entropy stability of the numerical scheme (i.e., entropy must only increase), which is generally not guaranteed (see also Section 5). Instead, in our proof-of-concept implementation, we simply turn off this implicit source term in regions with negative dissipation. We find that only a small fraction of cells is affected in our turbulence tests below, and the overall results remain the same.
Finally, it is worth noting that the above model does not affect the adiabatic evolution of gas. Indeed, only the fact that the residual energy ($e_{\rm tot}-e_{\rm kin}$) is partitioned between $e_{\rm th}$ and $e_{\rm turb}$ distinguishes it from ILES; however, given that both energies obey the same $\gamma=5/3$ equation of state, the pressure exerted on the gas remains exactly the same. This is also true when the dissipation term ($\epsilon$ in Equation (1)) is modeled explicitly, as it only changes the partitioning between $e_{\rm th}$ and $e_{\rm turb}$ but not their spatial distribution. This implies that the performance of the code in all idealized adiabatic tests without radiative cooling, like a shock tube or a point-source explosion, remains exactly the same as without this model. In the presence of cooling, on the other hand, the solution becomes different, as the model helps to retain some of the dissipated kinetic energy in the form of $e_{\rm turb}$, which decays on a timescale different from the thermal cooling timescale.
@@ 2.4 Turbulence Decay, Diffusion, and Extra Sources
Other than $\Sigma$, the rest of the explicit source terms in the second line of Equation (1) are the same in LES and SLES. Here we describe the adopted closures for these terms.
@@
The same term enters the equation for thermal energy (which is analogous to Equation (1)) but with a positive sign, implying that the dissipated subgrid turbulence appears as heat, which is then subject to radiative losses. The amount of dissipated subgrid turbulence, $\Delta e_{\rm turb}$, during a time step, $\tau$, is calculated from the analytic solution of $de_{\rm turb}/dt=-\epsilon$:
@@
with $\Delta e_{\rm turb}$ subtracted from $e_{\rm turb}$ and added to $e_{\rm th}$ after each step.
Next, for the turbulent diffusion term ($\boldsymbol{\nabla}\boldsymbol{\cdot}\kappa\boldsymbol{\nabla}e_{\rm turb}$), we adopt turbulent diffusivity of $\kappa\equiv C^{\prime}_{\kappa}\sigma\,\Delta=C_{\kappa}\sqrt{e_{\rm turb}/\rho}\Delta$ with $C_{\kappa}=\sqrt{2}\,C^{\prime}_{\kappa}=0.4$ (schmidt14; schmidt). A similar term is applied to all advected scalars, including thermal energy and chemical abundances, when the latter are followed explicitly. In our tests, the effect of explicit subgrid turbulent diffusion is generally smaller than numerical diffusivity except for the regions with very high $\sigma$ such as the interiors of SN bubbles in galaxy simulations (Section 4). The effect of turbulent diffusion is significantly more important in Lagrangian and quasi-Lagrangian (moving-mesh) codes owing to their reduced advection errors (e.g., rennehan21). As the choice of $C_{\kappa}$ does not affect our results, we do not consider it as a tunable parameter of the model, even though we keep the diffusion term on in the tests presented below. Turning off this term completely affects our results only weakly.
Finally, the explicit source term $S$ can be used to source subgrid turbulence via various unresolved processes, including stellar feedback. For example, a fraction of SN energy injected during a time step can be added directly to $e_{\rm turb}$ via this term. In our galaxy simulations presented in Section 4, we do not perform such an explicit injection (i.e., $S\equiv 0$), but we keep this term in Equation (1) for completeness.
@@ 3 Decaying supersonic turbulence
In this section, we use a decaying supersonic isothermal turbulence test to calibrate and investigate the performance of our new SLES model. To this end, we use a high-resolution simulation to predict the distribution of turbulent energy on small scales (Figure 2, top row). This distribution is then compared with low-resolution resimulations of the same initial conditions where the small-scale turbulence is predicted using the SLES model (Figure 2, bottom row). As we will show, despite its simplicity, the SLES model is able to capture multiple nontrivial features of the direct numerical simulation (DNS), such as the time evolution of the mean turbulent energy, its dependence on scale, and the correlation between gas density and turbulent velocity.
@@ 3.1 Numerical Setup and Analysis
Figure 2: Gas density (gray) and small-scale turbulent energy slices (orange) in the coarse-grained high-resolution simulation (DNS; the top set of panels) and low-resolution resimulations of the same ICs with the SLES model (the bottom set of panels). The maps are shown at $t=2/3\;t_{\rm cross,0}$. The leftmost panel in the top row shows the density slice at the native resolution of the DNS, while the other columns show coarse-grained values, with density being volume-weighted and $e_{\rm turb}$ computed from Equation (8); the corresponding grid size is indicated at the top. The bottom panels show slices at the same time step from low-resolution resimulations using the corresponding grid sizes, where $e_{\rm turb}$ is predicted using the SLES model. Thus, in total, four simulations are shown in the plot as indicated by the green shaded background: one DNS and three SLES resimulations. Low-resolution SLES resimulations produce smoother solutions than the DNS, but they capture the main global features of the small-scale turbulence, in particular, the correlation with gas density.
@@
To initialize the decaying isothermal turbulence test, we use the initial conditions from the code comparison project of kritsuk11. The initial sonic Mach number in this setup is $\mathcal{M}\approx 9$, and it decreases with time as turbulence is allowed to decay. The results below assume units in which the box size is $L_{\rm box}=1$, the average density is $\langle\rho\rangle=1$, and the constant sound speed is $c_{\rm s}=1$.
@@
After the start, the simulation undergoes a brief initial transient stage that lasts approximately one turbulent box-crossing time (see the slight bump in the dashed lines at $t<t_{\rm cross,0}$ in Figures 3 and 4 below). This transient is mainly caused by two factors. First, the initial conditions were created using a higher-order method (piecewise parabolic method) than we employ in our run. Second, the original simulation included magnetic pressure, which we removed in our runs. This pressure is, however, dynamically subdominant (corresponding to the Alfvén Mach number of $\approx 4.4$). Overall, the strength of the transient is rather small, and we present our analyses at sufficiently late times, $t_{\rm cross,0}$, after it is dissipated.
@@
As a reference for calibration and testing of the model, we run a high-resolution version of this setup with a $1024^{3}$ uniform grid. We do not model physical viscosity; instead, kinetic energy is dissipated numerically on scales close to the resolution. However, for comparisons with the SLES model, we compute the turbulent energy on scales larger than these numerical dissipation scales, on which the mean turbulent energy converges ($L\geq L_{\rm box}/64$; see Appendix A. For this reason, we henceforth refer to this reference simulation as a “direct numerical simulation” (DNS) even though, formally, the physical dissipation is absent in this simulation (the so-called “Coarse DNS” approach).
@@
The distribution of small-scale turbulence is derived by coarse-graining the direct simulation and averaging the properties of the flow on the scales below the coarse-graining scale. To this end, we use uniform $N^{3}$ grids, where $N<1024$ and is a power of 2 such that each coarse cell is a cubic volume with $1024/N$ cells on a side. To respect conservation laws, the averages of all conservative quantities are volume-weighted. For example, the coarse-grained density is computed as
@@
where $dV\equiv(L_{\rm box}/1024)^{3}$ and $dV_{N}\equiv(L_{\rm box}/N)^{3}$ are the cell volumes of the original $1024^{3}$ and coarse-grained $N^{3}$ grids.
@@
Implementation-wise, we perform the above averaging by exploiting the nested oct structure of the computational mesh in the ART code. Specifically, we use a static $16^{3}$ root grid where each cell is split into octs until the target uniform resolution of $1024^{3}$ is achieved, which corresponds to 7 refinement levels. Then, starting from the highest refinement level, hydrodynamic quantities are averaged between the cells in a given oct and saved level by level, providing a hierarchy of coarse-grained versions of the direct simulation.
The top row of Figure 2 shows the original high-resolution density map (the leftmost panel) and the series of coarse-grained versions using uniform grids with $N=128$, 64, and 32 cells on a side at $t=2/3\;t_{\rm cross,0}$, where $t_{\rm cross,0}$ is the initial turbulent box-crossing time. [^4] The specific time at which the maps are shown is picked such that the small-scale distribution of $e_{\rm turb}$ has time to develop, but the initial large-scale turbulent energy has not dissipated away. The top (gray) and bottom (orange) subsets of maps show, respectively, the coarse-grained density (Equation (6)) and the turbulent energy on the scales below the coarse-graining scale ($e_{\rm turb}$; Equation (8)). The distribution of $e_{\rm turb}$ shows a clear correlation with density: the higher turbulent energy corresponds to denser regions. In addition, the global mean value of $e_{\rm turb}$ increases with decreasing resolution of the coarse-graining grid, i.e., the map becomes more orange from left to right. This is because larger scales contain more turbulent energy.
@@
The results of such low-resolution resimulations with the SLES model are shown in the bottom set of panels. These low-resolution resimulations reproduce the overall features of the direct simulation, such as the locations of the high-density structures and the correspondence between high-density and high-$e_{\rm turb}$ regions. The solution is significantly smoother than in the DNS owing to higher numerical viscosity, i.e., larger numerical dissipation scales. The SLES model also captures the increase of the mean $e_{\rm turb}$ at a lower resolution.
Having such a set of simulations, the SLES model can be calibrated and tested by comparing its predictions with the coarse-grained DNS. The cell-by-cell comparison of $e_{\rm turb}$ is complicated due to the strong effect of resolution on the flow structure. Indeed, even minute changes due to, e.g., variations in the time-stepping scheme can cause slight offsets of the large-scale features like the locations of the shock fronts, leading to large cell-to-cell differences between the two runs (see, e.g., Figure 12 in gnedin18). For that reason, in the subsequent sections, we compare global statistics: the mean $e_{\rm turb}$ averaged over the entire box (which we also use to calibrate the model in Section 3.2); the normalization, slope, and scatter of the relation between gas density and $e_{\rm turb}$; and the power spectra of resolved kinetic energy, density, and $e_{\rm turb}$.
@@ 3.2 Model Calibration
Figure 3: Comparison of the global mean small-scale turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$, in coarse-grained DNS (dashed line) and low-resolution SLES resimulations with different values of the dissipation parameter, $C_{\epsilon}$ (colored lines). For this comparison, we use the grid size of 323, i.e., the cell size $\Delta\equiv L_{\rm box}/32$. In the DNS, $e_{\rm turb}$ is computed following Equation (8) while in the low-resolution resimulations, it is tracked by the SLES model. The mean $E_{\rm turb}$ is normalized to the total initial kinetic energy, i.e., $E_{\rm kin,0}\equiv E_{\rm turb}(<L_{\rm box})$ at $t=0$. The time axis is normalized by the initial box-crossing time of turbulence, $t_{\rm cross,0}\equiv 0.5\;L_{\rm box}/\sigma_{0}$, or $\approx 0.058$ sound-crossing times for the initial Mach number of $\mathcal{M}\approx 8.6$. The value of $C_{\epsilon}=1.6$ provides the best match to the global $E_{\rm turb}$ and therefore we will use it as our fiducial choice in the rest of the paper. The agreement with the DNS can be further improved by making $C_{\epsilon}$ dependent on the local subgrid Mach number, as shown with the thin line and described in the text.
The simplest version of the SLES model involves a single parameter that needs to be selected: $C_{\epsilon}$, describing the rate of subgrid turbulence dissipation in units of the local eddy cell-crossing time (Equation (3)). We calibrate the value of $C_{\epsilon}$ by matching the mean subgrid turbulent energy, $E_{\rm turb}\equiv\langle e_{\rm turb}\rangle_{\rm box}$.
@@
Note that as the model contains only a single parameter, only a single instantaneous value of $E_{\rm turb}$ from one of the time steps could be used for calibration. The fact that the SLES model also reasonably closely reproduces the temporal behavior of $E_{\rm turb}$ is nontrivial and can be viewed as a test of the SLES model rather than a part of calibration (see also Section 3.3.1). Indeed, as the global Mach number of decaying turbulence is decreasing, the values of $E_{\rm turb}$ separated by more than the global turbulence crossing time, $t_{\rm cross,0}$, can be interpreted as independent realizations of different Mach number regimes.
Finally, note that, although the simple SLES model captures the overall behavior of $E_{\rm turb}$ in the DNS, there are some differences in detail. In particular, around $t\sim 2\;t_{\rm cross,0}$, the value of $E_{\rm turb}$ in the SLES model overshoots DNS by $\sim 20\%$ and then decays at a somewhat faster rate, resulting in $\sim 20\%$ lower $E_{\rm turb}$ than that of DNS at $t\gtrsim 8\;t_{\rm cross,0}$.